# bayesNMF: Fast Bayesian Poisson NMF with Automatically Learned Rank Applied to Mutational Signatures

### v2

@@ 3.3 Baseline Models
The Poisson-Gamma model is available in SigneR [Rosales et al., 2016] while Truncated Normal and Exponential priors as well as Normal likelihoods are available in SignatureAnalyzer [Kasar et al., 2015, Kim et al., 2016]. We implement our own samplers to enable direct models comparisons where all settings but likelihoods and priors are held constant. We do not implement a standard Gibbs sampler for the Poisson-Truncated Normal model because the full conditionals are not from distributions that can be directly sampled, even with Poisson augmentation. The Poisson-Gamma model is the closest comparison.
To evaluate the ability of SBFI to learn rank, we compare it to minBIC, a procedure which fits a separate Poisson+MH sampler for each candidate rank and selects the one minimizing BIC. We also compare it to automatic relevance determination as implemented in SignatureAnalyzer, which learns rank but does not report posterior uncertainty.
Likelihood Prior Sampler Purpose Reference
Poisson Truncated Normal MH-within-Gibbs This Work -
Poisson Exponential MH-within-Gibbs This Work -
Poisson Gamma Standard Gibbs Baseline SigneR
Poisson Exponential Standard Gibbs Baseline SignatureAnalyzer
Normal Truncated Normal Standard Gibbs Motivation SignatureAnalyzer
Normal Exponential Standard Gibbs Motivation SignatureAnalyzer
Table 2: List of models implemented in bayesNMF. MH-within-Gibbs samplers are the new contribution of this work. Standard Poisson samplers serve as a baseline for comparison to validate MH-within-Gibbs models. Standard Normal samplers serve as motivation for the proposal distributions of MH-within-Gibbs models. References refer to other software implementing the same likelihood-prior combinations, though they are re-implemented in bayesNMF to allow for direct comparisons. (SignatureAnalyzer [Kasar et al., 2015, Kim et al., 2016], and SigneR [Rosales et al., 2016].)
@@ 3.4 Inference
Inference is performed on the final 1000 posterior samples. When learning rank, we first compute $\hat{A}$, the maximum a-posteriori (MAP) estimate of the signature inclusion matrix, at its mode. We then subset inference samples to those that match this MAP.
To summarize the inference samples, we next remove the scale indeterminacy of NMF by normalizing each iteration’s $P^{(i)}$ and $E^{(i)}$ matrices for columns of $P^{(i)}$ to sum to one. Then, final estimates $\hat{P}$ and $\hat{E}$ and 95% credible intervals are element-wise means and 2.5th and 97.5th percentiles, respectively. The estimated reconstruction $\hat{M}$ is the product $\hat{P}\hat{A}\hat{E}$.
To compare results with a reference $P$, we assign each inference samples’s signature matrix $P^{(i)}$ to $P$ with the Hungarian algorithm [Kuhn, 1955]. This assumes each reference signature may only be assigned to a single estimated signature, and vice versa. We ensemble assignments over inference samples with majority voting (details in Appendix B.4).
@@ 4 Implementation and R Software Package
4.1 Implementation
The bayesNMF R software package implements an automated convergence detection algorithm and reports convergence metrics in a log file and with trace plots (Figure 1A). This makes the software simple to use, reduces subjectivity of sampler stopping, and ensures all models reach a similar degree of convergence so results can be compared fairly. Full details of model specifications, Gibbs updates, and hyperparameters are available in Appendix A. Details of convergence, inference, and reference assignment are in Appendix B.
@@ 4.2 Modeling Capabilities
The bayesNMF R software package allows users to fit all models described in this paper with the bayesNMF function. Model specifications can be adjusted by the likelihood, prior, and MH parameters. The rank can be a fixed value or a range vector, in which case rank_method specifies whether minBIC, SBFI, or BFI is used to learn rank. Users are also able to set hyperprior parameters or specify initial values of parameters or prior parameters.
@@ 4.3 Reference Comparison and Visualization Capabilities
Figure 2: Relative performance of Bayesian NMF models. Grey fill represents Gamma prior on standard Poisson models, but Truncated Normal prior on Poisson+MH or Normal models. White fill always indicates Exponential priors. A. Agreement of $\hat{E}$ (left) and $\hat{P}$ (right) between models as signature-wise minimum cosine similarities. Top: standard Poisson vs Normal; Middle: standard Poisson vs Poisson+MH; Bottom: standard Poisson vs standard Poisson (null case). Excluding 0 to 7 values per density for similarity <<0.95 for visual clarity (see Appendix Table C.4). B. Efficiency gain of Poisson+MH relative to standard Poisson. C. Performance of novel Poisson+MH models.
@@ 5.1 Metrics
Before computing metrics, we align signatures as described in Section 3.4. SignatureAnalyzer does not report posterior samples so results are aligned on final estimates only.
When rank is known and two models are compared to one another, we report the cosine similarity of the worst-aligned signature, as well as that of the worst-aligned exposure profile. If the two solutions are identical, both values will be exactly 1. We also report relative efficiency as the ratio of total runtime between two models.
When rank is known and a model is compared to ground truth, we first evaluate NMF solutions in terms of their ability to reconstruct the original data matrix using RMSE and KL-Divergence between $M$ and $\hat{M}$. We then measure their ability to reconstruct ground truth latent factors using the cosine similarity of the worst-reconstructed signature. Finally, we report total runtime in minutes to assess computational efficiency.
When learning rank, we evaluate NMF solutions by the bias of estimated rank as well as post-alignment sensitivity and precision. Sensitivity is the proportion of true signatures that have an estimated match of cosine similarity >>0.9. It is penalized by underestimating rank because at least one true signature will be without an estimated match. Precision is the proportion of estimated signatures that have a true match. It is penalized by overestimating rank. It may also be penalized by underestimating rank if estimated signatures are combinations of true signatures, but not if one or more true signatures are simply excluded.
@@ 5.2 Poisson+MH efficiently samples the correct posterior
In this section, we further justify our Normal model-based proposals, and we show that Poisson+MH models more efficiently sample the same posterior as standard Poisson models.
We consider true latent ranks $N\in[2,4,8,16]$ and sample sizes $G\in[16,32,64,128]$ such that $G\geq 8N$. For each combination, 10 datasets are generated. $N$ signatures are sampled from COSMIC SBS v3.3.1 [Tate et al., 2019], each normalized to sum to 1. The expected number of mutations for a sample $g$ is $m_{g}\sim\text{Negative Binomial}(r\cdot N,p)$}, with $r=111.11$ and $p=0.1$. This yields $E[m_{g}]=1000\cdot N$, or $1000$ mutations per signature per sample, on average, so that each simulation provides comparable relative power to learn each signature. Columns of the exposures matrix, $E$, are generated $E_{g}\sim\text{Multinomial}(m_{g},\mathbf{p})$, where $\mathbf{p}\sim\text{Dirichlet}(\mathbf{1}_{N})$. Mutational counts are simulated as $M_{kg}\sim\text{Poisson}(\lambda_{kg}=(PE)_{kg})$.
Figure 2A compares point estimates. In all comparisons, agreement is >>0.99 in at least 90% of cases for $\hat{E}$ and 84% for $\hat{P}$. The top row shows Normal Bayesian NMF results in point estimates nearly identical to Poisson Bayesian NMF, justifying our Normal model-based MH proposals. The middle row shows the addition of MH steps does not alter point estimates of Poisson Bayesian NMF. The bottom row is a null case, comparing two standard Poisson samplers, to show that some variation is expected due to the randomness of Gibbs samplers. Along with aligned point estimates, the addition of MH steps increases effective sample sizes (Appendix C.1) and does not substantially change credible interval widths (Appendix C.3). Together, these results show the correctness of our novel Poisson+MH samplers in that they explore the same posterior space as the standard Poisson samplers.
Figure 2B reports relative efficiency of Poisson+MH compared to standard Poisson samplers. Poisson+MH models are 3-30x faster, and the gap grows with dimensionality.
Figure 2C reports metrics of the Poisson+MH samplers as a baseline for the comparisons of Figures 2A-B. Metrics for all models are available in Appendix Figure C.1.
@@ 5.3 SBFI learns rank
In this section, we demonstrate the utility of SBFI for automatically learning rank. SBFI performs comparably to, or better than, SignatureAnalyzer’s automatic relevance determination (ARD) and reports posterior uncertainty. Our recommendation is bayesNMF with SBFI for Truncated Normal priors or bayesNMF with minBIC for Exponential priors.
@@
Data are simulated as in Section 5.2, but with $G=64$ and true ranks $N$ between $1$ and $20$. 10 datasets are simulated per $N$ such that no two included signatures have cosine similarity $\geq$0.8. We note that lower precision and sensitivity at the larger ranks may reflect the difficulty of the problem rather than flaws of the methods considered.
Figure 3 reports rank bias, precision, sensitivity, and runtime for each model as true rank increases along the x-axis. For a Truncated Normal / L2 prior (filled shapes), all approaches estimate rank very well, but with different error patterns. SignatureAnalyzer overestimates low and high ranks, while SBFI and minBIC both underestimate high ranks. At high ranks, SBFI and minBIC perform best in terms of precision and sensitivity, meaning that the signatures our models do estimate are correct, even though some are missed, while SignatureAnalyzer yields some noisy or split signatures even if it learns rank correctly. For an Exponential / L1 prior (empty shapes), the sparsity penalty of SBFI may be too strong, as SBFI underestimates high ranks by a wide margin (up to 8 signatures under for $N=20$). However, minBIC still outperforms SignatureAnalyzer in terms of precision and sensitivity.
@@ 6.1 Data
In this section, we analyze the Pan-Cancer Analysis of Whole Genomes (PCAWG) mutational counts data for whole genome sequencing (WGS) 96-alphabet single base substitution (SBS) mutation classification Consortium [2020]. This is publicly available from the International Cancer Genome Consortium’s (ICGC) Accelerating Research in Genomic Oncology (ARGO) data portal [Zhang et al., 2019, access link]. This includes 37 datasets, each for a different histology group. We excluded histology groups with 10 or fewer samples, as well as benign (non-cancerous) histology groups, leaving 32 datasets for analysis.
We exclude hypermutated samples from all analyses. The inclusion of hypermutated samples can result in signature bleeding, where mutations in non-hypermutated samples are misattributed to hypermutation-specific signatures [Maura et al., 2019, Alexandrov et al., 2020]. For each histology group, we identify hypermutated samples by assuming the total mutation count per sample follows a Negative Binomial mixture model, with at least one component for non-hypermutated samples, and possibly one or more components for hypermutated samples (details in Appendix D.1). This removes 65 samples total (<<3% of data), with at most 13 samples removed per histology group in Skin-Melanoma.
@@ 6.2 Methods
Because ground truth is unknown for real data applications, rigorous conclusions can only be made in the context of simulated data as in Section 5. However, we wish to illustrate our method and compare it to others in a real world setting. We compare our results to SignatureAnalyzer and show our approach results in similar biological interpretations.
We use Poisson-Truncated Normal+MH SBFI as this specification is the closest to the Gamma priors used in most mutational signatures literature. We fit a separate sampler to each histology group, learning ranks $\in[1,20]$. For comparison, we also fit SignatureAnalyzer with a Poisson objective and L2 priors (equivalent to Truncated Normal priors). Estimated signatures are assigned to the COSMIC reference as described in Section 3.4.
Figure 4: Results of bayesNMF+SBFI (upwards triangle) and SignatureAnalyzer+ARD (downwards triangle) on PCAWG histology groups. Skin Melanoma is excluded for visual clarity because SignatureAnalyzer estimates a high rank of 76, (bayesNMF estimates a rank of 10). A) COSMIC reference signatures aligned to $\hat{P}$. Reference signatures not found with either method are excluded. B) Latent ranks estimated by bayesNMF+SBFI and SignatureAnalyzer+ARD. Exact values in Appendix Table D.5.
@@ 6.3 Results
Figure 4A shows a high degree of overlap in the signatures discovered by bayesNMF+SBFI and SignatureAnalyzer+ARD. These results mirror many trends from Alexandrov et al. [2020], including the presence of SBS1, SBS5, and SBS40 in nearly every cancer type. More detailed results including the median number of mutations attributed to each signature in each histology group by each method can be found in Appendix Figure D.2.
A clear difference between methods is the estimated rank—bayesNMF+SBFI typically results in a sparser solution, with SignatureAnalyzer+ARD frequently estimating twice the number of signatures (Figure 4B). Based on our simulation results, because bayesNMF+SBFI with Truncated Normal priors consistently achieved higher sensitivity and precision, we speculate that results here of SignatureAnalyzer with L2 priors may include spurious components (especially for Skin Melanoma, where SignatureAnalyzer estimates a rank of 76), making the resulting signatures less biologically interpretable.
In most cases where bayesNMF recovered a signature not reported by SignatureAnalyzer (Figure 4A upwards triangles), the discrepancy can be partly explained by posterior uncertainty. These signatures have lower cosine similarity with their assigned reference signatures and a lower proportion of assignment votes, indicating uncertainty in assignment (Appendix Figure D.3). However, of the 34 signatures of this type, 28 (82%) had a SignatureAnalyzer-derived signature among the candidates receiving assignment votes, just not the majority required for posterior ensemble assignment. Therefore, the signature estimated by SignatureAnalyzer can be viewed as consistent with one plausible posterior mode inferred by bayesNMF. There are 6 signatures still unique to bayesNMF after considering this aspect, of which many match previous discoveries in the COSMIC database: SBS5 in CNS pilocytic astrocytoma and pancreatic adenocarcinoma, SBS1 in lymphocytic leukemia (CLL), and SBS95 in liver hepatocellular carcinoma (HCC). Some are not in the COSMIC database, but have been previously reported in the literature: SBS36 in prostate adenocarcinoma [Jiang et al., 2020]. However, there is no reference for SBS9 in esophageal adenocarcinoma, although it has been found in esophageal small cell carcinoma.

### v3

@@ 3.3 Baseline Models
The Poisson-Gamma model is available in SigneR [Rosales et al., 2016] while Truncated Normal and Exponential priors as well as Normal likelihoods are available in SignatureAnalyzer [Kasar et al., 2015, Kim et al., 2016]. We implement our own samplers to enable direct model comparisons where all settings but likelihoods and priors are held constant. We do not implement a standard Gibbs sampler for the Poisson-Truncated Normal model because the full conditionals are not from distributions that can be directly sampled, even with Poisson augmentation. The Poisson-Gamma model is the closest comparison.
To evaluate the ability of SBFI to learn rank, we compare it to minBIC, a procedure that fits a separate Poisson+MH sampler for each candidate rank and selects the one minimizing BIC. We also compare it to automatic relevance determination as implemented in SignatureAnalyzer, which learns rank but does not report posterior uncertainty. Under the non-negativity constraints in NMF, SignatureAnalyzer’s L1 and L2 priors correspond to Exponential and Truncated Normal priors, respectively.
Likelihood Prior Sampler Purpose
Poisson Truncated Normal MH-within-Gibbs This Work
Poisson Exponential MH-within-Gibbs This Work
Poisson Gamma Standard Gibbs Baseline
Poisson Exponential Standard Gibbs Baseline
Normal Truncated Normal Standard Gibbs Motivation
Normal Exponential Standard Gibbs Motivation
Table 2: List of models implemented in bayesNMF. MH-within-Gibbs samplers are the new contribution of this work. Standard Poisson samplers serve as a baseline for comparison to validate MH-within-Gibbs models. Standard Normal samplers serve as motivation for the proposal distributions of MH-within-Gibbs models.
@@ 3.4 Inference
For results in Sections 5 and 6, inference is performed on the final 1000 posterior samples, though our software allows this number to be user-specified and for all samples to be stored.
To summarize the posterior distribution when rank is fixed, we first remove the scale indeterminacy of NMF by normalizing each iteration’s $P^{(i)}$ and $E^{(i)}$ matrices so that columns of $P^{(i)}$ sum to one. Then, posterior means $\hat{P}$ and $\hat{E}$ are element-wise averages, and 95% credible intervals are element-wise 2.5th and 97.5th percentiles.
Parameters $P$ and $E$ are only interpretable in the context of the inclusion structure $A$. Therefore, when learning rank, the target for inference is the posterior distribution conditional on the maximum a-posteriori (MAP) inclusion matrix $\hat{A}$, identified as the mode across samples $A^{(i)}$. Summaries of the conditional posterior are then computed from samples where $A^{(i)}=\hat{A}$. Within the selected samples, $\hat{P}$ and $\hat{E}$ are computed as above.
For optional post-hoc comparison with a reference $P$, we assign each $P^{(i)}$ to $P$ with the Hungarian algorithm [Kuhn, 1955]. This assumes each reference signature may only be assigned to a single estimated signature, and vice versa. We ensemble assignments across samples with majority voting (details in Appendix B.4).
@@ 4 Implementation and R Software Package
4.1 Software
The bayesNMF R software package is available for download on GitHub at jennalandy/bayesNMF. Details of model specifications, Gibbs updates, and hyperparameters are available in Appendix A. Details of automated convergence detection and diagnostics, inference, and reference assignment are in Appendix B.
@@ 4.2 Modeling Capabilities
Users are able to fit all models described in this paper with the bayesNMF function. Model specifications can be adjusted by the likelihood, prior, and MH parameters. The rank can be a fixed value or a range vector, in which case rank_method specifies whether minBIC, SBFI, or BFI is used to learn rank. Users are also able to set hyperprior parameters or specify initial values of parameters or prior parameters.
@@ 4.3 Reference Comparison and Visualization Capabilities
Figure 2: Relative performance of Bayesian NMF models. Grey fill represents Gamma priors on standard Poisson models, but Truncated Normal priors on Poisson+MH or Normal models. White fill always indicates Exponential priors. A. Agreement of $\hat{E}$ (left) and $\hat{P}$ (right) between models as signature-wise minimum cosine similarities. Excluding 0 to 7 values per density for similarity <<0.95 for visual clarity (see Appendix Table C.1). B. Efficiency gain of Poisson+MH relative to standard Poisson. C. Performance of novel Poisson+MH models. Metrics for all models are available in Appendix Figure C.1.
@@ 5.1 Metrics
Before computing metrics, signatures are aligned to ground truth. Results for bayesNMF are aligned as described in Section 3.4. SignatureAnalyzer does not report posterior samples, so its results are aligned using point estimates only.
When rank is known, and two models are compared to one another, we report the cosine similarity of the worst-aligned signature, as well as that of the worst-aligned exposure profile. If the two solutions are identical, both values will be exactly 1. We also report relative efficiency as the ratio of total runtime between two models.
When rank is known, and a model is compared to ground truth, we first evaluate NMF solutions in terms of their ability to reconstruct the original data matrix using RMSE and KL-Divergence between $M$ and $\hat{M}=\hat{P}\hat{E}$. We then measure their ability to reconstruct ground truth latent factors using the cosine similarity of the worst-reconstructed signature. Finally, we report the total runtime in minutes to assess computational efficiency.
When learning rank, we evaluate NMF solutions by the bias of the estimated rank as well as post-alignment sensitivity and precision. Sensitivity is the proportion of true signatures that have an estimated match of cosine similarity >>0.9. It is penalized by underestimating rank because at least one true signature will be without an estimated match. Precision is the proportion of estimated signatures that have a true match. It is penalized by overestimating rank. It may also be penalized by underestimating rank if estimated signatures are combinations of true signatures, but not if one or more true signatures are simply excluded.
@@ 5.2 Poisson+MH efficiently samples the correct posterior
In this section, we treat rank as known and perform model comparisons to further justify our Normal model-based proposals and to show that Poisson+MH models more efficiently sample the same posterior as standard Poisson models.
We consider true latent ranks $N\in[2,4,8,16]$ and sample sizes $G\in[16,32,64,128]$. For each $N,G$ where $G\geq 8N$, 10 datasets are generated. Signatures are sampled from COSMIC SBS v3.3.1 [Tate et al., 2019], each normalized to sum to 1. The expected total mutation count for sample $g$ is $m_{g}\sim\text{Negative Binomial}(111.11N,0.1)$}, yielding $E[m_{g}]=1000N$, or $1000$ mutations per signature per sample. This gives each simulation comparable relative power to learn each signature, and this explains why reconstruction errors increase with $N$. Columns of the exposures matrix, $E$, are generated $E_{g}\sim\text{Multinomial}(m_{g},\mathbf{p}_{g})$, where $\mathbf{p}_{g}\sim\text{Dirichlet}(\mathbf{1}_{N})$. Mutation counts are simulated as $M_{kg}\sim\text{Poisson}((PE)_{kg})$.
Normal Bayesian NMF results in point estimates nearly identical to Poisson Bayesian NMF, justifying our Normal model-based MH proposals (Figure 2A, top row). Further, the addition of MH steps does not alter point estimates of Poisson Bayesian NMF (middle row). These comparisons yield agreement >>0.99 in at least 84% of cases and are very close to the null case comparing two independent chains of the standard Poisson sampler (bottom row). Along with aligned point estimates, the addition of MH steps increases effective sample sizes (Appendix C.1) and does not substantially change credible interval widths (Appendix C.3). Together, these results show that our novel Poisson+MH samplers yield the same posterior inference as the standard Poisson samplers. Additionally, Poisson+MH models are 3-30x faster, and this gap grows with dimensionality (Figure 2B).
@@ 5.3 SBFI learns rank
In this section, we demonstrate that SBFI performs comparably to or better than SignatureAnalyzer’s automatic relevance determination (ARD) while reporting posterior uncertainty.
@@
Data are simulated as in Section 5.2 with $G=64$ and 10 datasets for each true rank $N$ between $1$ and $20$. For each dataset, no two signatures have a cosine similarity $\geq$0.8. Reduced precision and sensitivity observed at larger ranks likely reflect the difficulty of the problem (i.e., large $N$ relative to $G$), rather than deficiencies of the methods considered.
Figure 3 reports rank bias, precision, sensitivity, and runtime for each model as true rank increases along the x-axis. For a Truncated Normal / L2 prior (filled shapes), all approaches estimate rank well, but with different error patterns: SignatureAnalyzer overestimates low and high ranks, while SBFI and minBIC underestimate high ranks. Under this prior and at high ranks, SBFI and minBIC achieve the highest precision and sensitivity, meaning that the signatures our models do estimate are correct, even though some are missed, while SignatureAnalyzer often yields noisy or split signatures, even if it learns the correct rank. For an Exponential / L1 prior (empty shapes), the sparsity penalty of SBFI may be too strong, as SBFI underestimates high ranks by a wide margin. In this setting, minBIC still shows higher precision and sensitivity than SignatureAnalyzer.
@@ 6.1 Data
In this section, we analyze the Pan-Cancer Analysis of Whole Genomes (PCAWG) mutation count data using the 96-alphabet single base substitution (SBS) mutation classification [Consortium, 2020]. This is publicly available from the International Cancer Genome Consortium’s (ICGC) Accelerating Research in Genomic Oncology (ARGO) data portal [Zhang et al., 2019]. This includes 37 datasets, each for a different histology group. We excluded histology groups with 10 or fewer samples, as well as benign (non-cancerous) histology groups, leaving 32 datasets for analysis.
We exclude hypermutated samples from all analyses. The inclusion of hypermutated samples can result in signature bleeding, where mutations in non-hypermutated samples are misattributed to hypermutation-specific signatures [Maura et al., 2019, Alexandrov et al., 2020]. For each histology group, we identify hypermutated samples by assuming the total mutation count per sample follows a Negative Binomial mixture model, with at least one component for non-hypermutated samples, and possibly one or more components for hypermutated samples (details in Appendix D.1). This removes 65 samples total (<<3% of data), with at most 13 samples removed per histology group in skin melanoma.
@@ 6.2 Methods
Because ground truth is unknown for real data applications, rigorous conclusions can only be made in the context of simulated data, as in Section 5. However, we wish to illustrate our method and compare it to others in a real-world setting. We compare our results to SignatureAnalyzer and show that our approach results in similar biological interpretations.
We use the bayesNMF Poisson+MH approach with Truncated Normal priors and SBFI for rank $\in[1,20]$. For each histology group, we fit bayesNMF+MH, and compare to SignatureAnalyzer with L2 priors.
Figure 4: Results of bayesNMF+SBFI (upwards triangle) and SignatureAnalyzer+ARD (downwards triangle) on PCAWG histology groups. Skin melanoma is excluded for visual clarity because SignatureAnalyzer estimates a high rank of 76 (bayesNMF estimates a rank of 10). A. COSMIC reference signatures aligned to $\hat{P}$. Reference signatures not found with either method are excluded. B. Estimated latent ranks. Exact values in Appendix Table D.1.
@@ 6.3 Results
Figure 4A shows a high degree of overlap in the signatures discovered by bayesNMF+SBFI and SignatureAnalyzer+ARD. These results mirror many trends from Alexandrov et al. [2020], including the presence of SBS1, SBS5, and SBS40 in nearly every cancer type. More detailed results, including the median number of mutations attributed to each signature in each histology group by each method, can be found in Appendix Figure D.2.
A clear difference between methods is the estimated rank—bayesNMF+SBFI typically results in a sparser solution, with SignatureAnalyzer+ARD frequently estimating twice the number of signatures (Figure 4B). Based on our simulation results, because bayesNMF+SBFI with Truncated Normal priors consistently achieved higher sensitivity and precision, we speculate that the results here of SignatureAnalyzer with L2 priors may include spurious components (especially for skin melanoma, where SignatureAnalyzer estimates a rank of 76), making the resulting signatures less biologically interpretable.
In most cases where bayesNMF recovered a signature not reported by SignatureAnalyzer (Figure 4A upwards triangles), the discrepancy can be partly explained by posterior uncertainty. These signatures have lower cosine similarity with their assigned reference signatures and a lower proportion of assignment votes, indicating uncertainty in assignment (Appendix Figure D.3). However, of the 34 signatures of this type, 28 (82%) had a SignatureAnalyzer-derived signature among the candidates receiving assignment votes, just not the majority required for posterior ensemble assignment. Therefore, many signatures estimated by SignatureAnalyzer can be viewed as consistent with one plausible posterior mode inferred by bayesNMF. Still, there are 6 signatures unique to bayesNMF after considering this aspect, of which many match previous discoveries in the COSMIC database: SBS5 in CNS pilocytic astrocytoma and pancreatic adenocarcinoma, SBS1 in lymphocytic leukemia (CLL), and SBS95 in liver hepatocellular carcinoma (HCC). Some are not in the COSMIC database, but have been previously reported in the literature: SBS36 in prostate adenocarcinoma [Jiang et al., 2020]. There is no reference for SBS9 in esophageal adenocarcinoma, although it has been found in esophageal small cell carcinoma.